Of 85 randomly selected women, 38 have tried some form of alternative medicine. Of 90 randomly selected men, 23 have tried some form of alternative medicine. At the 1% significance level, do the data provide sufficient evidence to conclude that women are more likely than men to try alternative forms of medicine? *Please show all work and do both the P val approach and crit value thanks
Answer #1

H0 : p1 = p2
HA: p1 > p2

test statistics:

p1 = 38/85 = 0.447
p2 = 23/90 = 0.256
pcap = 38 + 23/85 + 90 = 0.3486

z = ( p1 - p2) /sqrt(pcap*(1-pcap) *(1/1n+1/n2))
= ( 0.447 - 0.256)/sqrt(0.3486 *(1-0.3486)*(1/85 + 1/90)
= 2.6571

p value = .0039

Critical value = 2.33

There is sufficient evidence to conclude that women are more likely than men to try alternative forms of medicine

Reject H0